OpenStudy (anonymous):

if f(x)= h(x)g(x), where g(0)=2, g'(0)=5, h(0)=3, and h'(0)=-9, find f'(0) (step by step)

OpenStudy (freckles):

Use the product rule since you have a product. (uv)'=uv'+vu'

OpenStudy (freckles):

After you have applied product rule you can then replace all the x's with 0 and use the values given to you to find f'(0)

OpenStudy (anonymous):

am i supposed to do this........h(x)g(x)'+g(x)h(x)' @freckles

OpenStudy (freckles):

\[f'(x)=g'(x)h(x)+g(x)h'(x)\]

OpenStudy (freckles):

Usually people put the derivative sign thing directly behind the function name

OpenStudy (anonymous):

do i plug in the zero's or the values into this function? @freckles

OpenStudy (freckles):

You want to find f'(0) and f'(x)=g'(x)h(x)+g(x)h'(x) so what do you think we need to replace x with?

OpenStudy (anonymous):

zero

OpenStudy (freckles):

yes so do so

OpenStudy (anonymous):

what takes place of the letters? @freckles

OpenStudy (freckles):

you mean what is g(0) and g'(0) and h(0) and h'(0)?

OpenStudy (freckles):

These values are all given to you in your question.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

am i supposed to do this...... f(o)=5(0)93(0))+2(0)(-9(0)) @freckles

OpenStudy (freckles):

woah.. wait...did you replace all the x's with 0's in the equation: f'(x)=g'(x)h(x)+g(x)h'(x)

OpenStudy (freckles):

f'(0)=g'(0)h(0)+g(0)h'(0)

OpenStudy (freckles):

look at your question and look what is given to you

OpenStudy (freckles):

you are given g'(0) is 5 so replace g'(0) with 5 you are given h(0) is 3 so replace h(0) with 3 you are given g(0) is 2 so replace g(0) with 2 you are given h'(0) is -9 so replace h'(0) with -9

OpenStudy (anonymous):

f'(0)=5(3)+2(-9) f'(0)= -3 @freckles

OpenStudy (freckles):

exactly

OpenStudy (anonymous):

thank you! @freckles